A rectangular parcel of land is 20 m wide. The length of the diagonal between the opposite corners is 5 m more than the length of the parcel. What is the length of the parcel?
Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

A rectangular parcel of land is 20 m wide. The length of the diagonal between the opposite corners is 5 m more than the length of the parcel. What is the length of the parcel?
---
Draw the picture of the rectangle.
width = 20 m
length = x meters
diagonal = x+5 meters
----
Solve for "x"::
(x+5)^2 = x^2 + 400
----
x^2+10x + 25 = x^2 + 400
10x = 375
x = 37.5 meters (length)
------------
Cheers,
Stan H.
----------

Answer by Cromlix(4381)

Hi there,
Imagine you have a right angled triangle.
Width = 20 meters
Length = x
Diagonal = x + 5
(20)^2 + (x)^2 = (x + 5)^2
400 + x^2 = x^2 + 10x + 25
Collect like terms.
10x = 400 - 25
10x = 375
x = 37.5 meters
Length of parcel = 37.5 meters.
Hope this helps :-)